# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v1

@@ 3.0.2 Knowledge Distillation Methods
Caron et al. (2021) introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
Li et al. (2022) proposed Efficient Self-Supervised ViTs (EsViT) to achieve a optimized balance between accuracy and efficiency, reducing the cost in building state-of-the-art SSL vision systems while demonstrating enhanced scaling performance on accuracy versus throughput and model size. They introduced a multi-stage Transformer architecture, combining networks to address the issue of semantic information loss when transitioning from a single-stage transformer to a multi-stage architecture. Employing a distillation learning strategy targeting correspondence between patches, they introduced a new non-contrastive region-matching pretraining task, enhancing the model’s ability to learn intricate regional dependencies and substantially improve the quality of visual representations learned.
@@
Xie et al. (2021) introduced MoBY, a SSL framework tailored for ViTs. This approach combines MoCo v2 and BYOL approaches, yielding notably high accuracy on the ImageNet-1K dataset. MoBY adopts the key queue, momentum design, and contrastive loss from MoCo v2, and it also take asymmetric data augmentations, asymmetric encoders, and momentum scheduler from BYOL. Employing a knowledge distillation approach, MoBY employees a pair of neural networks (online and target) to enhance mutual learning. Both encoders includes a backbone and a projector head, with the online encoder featuring an additional prediction head, rendering the two encoders non-uniform. Gradients progressively optimizes the online encoder, while the target encoder is a moving average of the online encoder updated through momentum. MoBY is evaluated across diverse downstream tasks, such as object detection and semantic segmentation, showcasing robustness and transferability across varied visual tasks. When utilizing DeiT-S and Swin-Transformers as backbone architectures, MoBY achieves top-1 accuracy rates of 72.8% and 75.0%, respectively, on ImageNet-1K.
@@ 3.0.3 Clustering Methods
Su et al. (2023) introduced a clustering-based SSL framework tailored for ViTs, aiming to enhance performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view which considers over an entire dataset. FLSL utilizes a bi-level feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally.
@@ 3.0.4 Generative Methods
123456789- Chen et al. (2020) introduced iGPT, an image-based version of GPT, for self-supervised visual learning. Unlike ViT, which uses patch embedding, iGPT directly resizes and flattens images into lower-resolution sequences, which are then fed into a GPT-2 model for autoregressive pixel prediction. Despite its significant computational cost, iGPT achieves high accuracy on CIFAR-10, surpassing supervised Wide ResNet.
Bao et al. (2022) introduced BEIT, that is Bidirectional Encoder representation from Image Transformers is a self-supervised vision representation model. Instead of pixel-wise generation, they adopted MIM as a task to pretrain ViTs in a self-supervised manner. BEIT, grounded on a BERT-style visual Transformer, reconstructs masked images in the latent space. The approach involves randomly masking parts of image pieces and predicting the visual tokens relevant to these hidden pieces. The main motivation behind this pre-training technique is to restore the base visual tokens utilizing corrupted image patches. The standard visual Transformer serves as the backbone network. Evaluation using thorough fine-tuning for tasks such as image classification and semantic segmentation demonstrated competitive performance without human annotation.
He et al., (2021) proposed MAE (Masked Autoencoders), in this methodology parts of input images are randomly masked and subsequently reconstructed the pixels that were masked, enabling optimized training of large models that demonstrate strong generalize across downstream tasks. This method utilizes an asymmetric encoder-decoder setup, where the encoder only looks on the visible parts of image, and a small decoder rebuilds the original image usingthe learned representation and mask tokens. By hiding a large proportion of the input image, e.g., 75%, MAE creates a challenging and useful self-supervised learning task. This scalable method allows large models to train efficiently and effectively, resulting in powerful models that generalize well. For instance, a standard ViT-Huge model achieved the highest accuracy of (87.8%) among techniques using only ImageNet-1K data. Supervised pretraining was outperformed by transfer learning in downstream task, showing strong potential for scaling behavior.
@@ 3.0.5 Hybrid Methods
Quan et al. (2023) introduced a novel SSL model, based on the global contrast-masked autoencoder (GCMAE), which adeptly extracts both global and local features from pathological images. GCMAE leverages masking image reconstruction and contrastive learning as self-supervised pretext tasks, enabling the encoder to proficiently depict local-global features. This model comprises of an initializer, an encoder, a tile feature extractor, and a global feature extractor facilitating image reconstruction and contrastive learning tasks. In this investigation, an asymmetric encoder-decoder architecture is employed. ViT serves as the backbone for the encoder, pretrained using MAE. Its module related to patches feature extraction is comprised of eight transformer blocks, forming an asymmetric structure in conjunction with an encoder featuring of about 12 transformer blocks (ViT-base). This asymmetry facilitates decoupling of the encoder and decoder, fostering the acquisition of more generalized representations. Extraction of global feature is achieved via utilizing contrastive learning. The hidden representation Vpatch(vis) is used to reconstruct image through decoder and update a memory bank (B). This updation of B is executed to store the global features utilizing momentum coefficient t. The dynamically update is conducted to pursue the feature vectors of image data using memory bank as a fixed size queue. The various features that are randomly picked from the memory bank are used as a negative samples that helps to reduce the effects of the constraint caused by batch size on the heavy performance. Notably, the momentum adjustment mechanism is designed distinctively from MoCo’s model parameter update, it employ an encoder and a memory bank to facilitate contrastive learning. Vpath(vis) is depicted as the encoders output in the current epoch in the latent representation, where as the vpatch(vis) represented the previous epoch latent representation stored in B. Particularly, the latent demonstrations at the output of encoder as vpatch(vis) in the current epoch, The cost function combines both, the weighted sum of MSE loss for tile feature extraction and the NCE loss for global feature extraction. This formulation aims to minimize the gap between similar features, enhancing the model’s generalization and accuracy, particularly in cross-dataset transfer learning tasks.
@@
Fang et al. (2023) proposed a hybrid SSL framework (HSL) for Hyperspectral Image (HSI) classification. Leveraging MIM and contrastive learning to enhance its effectiveness. Employing a ViT as the backbone network within an asymmetric encoder-decoder architecture with two branches, HSL achieve improved HSI classification utilizing contrastive pre-training and self-supervised generative tasks by effectively extracting spatial and spectral details. To facilitate image reconstruction through masked inputs, each branch connects intermediate features of the encoder to the decoder at the corresponding level. Contrastive learning is accomplished by utilizing the intermediate features that are obtained from two encoders that share parameters across both branches. Utilizing standard data augmentation methods, HSL generates several views of input images. Then each image view undergoes random masking technique such as MAE. Initially, image segments with mask are reorganised into a series of patches and later processed by HSL encoder and decoder to extract patterns. Extracted features are then pass through an adaptive average pooling layer and a projection head to achieve the contrastive learning task of capturing image resemblance. The decoder, which also includes multiple Transformer encoders, rebuilds patches after certain patches have been masked. Moreover, the Transformer Encoder retains the initial structure while the two output features from the encoder are used to compute contrastive loss. The loss is calculated using the decoded sequence produced by the decoder and the sequence that precedes the encoder masking, efficiently completing the masked image rebulding process. In summary, HSL combines contrastive pre-training and self-supervised generative tasks, facilitating effectiveness for downstream HSI classification task.
@@ 4 Regularization Techniques
Bootstrap Your Own Latent (BYOL) employs a teacher-student framework, where the student network predicts the output of a teacher network on augmented inputs. The teacher network is updated as a moving average of the student’s parameters, ensuring learning consistency. This method mirrors the stabilization effect seen in regularization techniques like batch normalization and parameter sharing.
@@
Momentum Contrast (MoCo) leverages a momentum-based moving average of the query encoder, establishing a large and consistent dictionary. This approach acts as a dynamic regularization mechanism, ensuring stability in the representation of dictionary keys over time. This stability is vital for the extraction of stable and generalizable features, akin to how traditional regularization methods prevent overfitting by maintaining consistency in model parameters.
@@
SimCLR utilizes a straightforward contrastive learning framework, where its regularization is primarily through aggressive data augmentation techniques. These augmentations include random cropping, resizing, color distortion, and Gaussian blur, enabling the model to learn invariant and robust features. This is similar to how techniques like dropout and data augmentation in traditional settings prevent over-reliance on specific input features.
@@
DINO (Self-Distillation with No Labels) utilizes self-distillation for regularization, where the student network mimics the teacher network, updated as an exponential moving average of the student’s parameters. This method is reminiscent of techniques like early stopping and weight decay, which aim to prevent overfitting by smoothing the learning trajectory.
@@
Barlow Twins method aims to align the cross-correlation matrix between the outputs of twin networks with the identity matrix. This alignment encourages the model to learn features invariant to input distortions, functioning as a regularization technique by discouraging the learning of redundant features.
@@
SwAV (Swapping Assignments between Views) regularization is unique, involving clustering approaches and consistency enforcement between cluster assignments for different views of the same image. This approach is akin to enforcing a constraint (regularization) on the feature space to enhance generalizability.
@@
TiCo (Time-Contrastive Consistent Learning of Visual Representations) focuses on learning temporally consistent representations, acting as a regularization mechanism by ensuring robustness to temporal variations in data. This is similar to entropy regularization, which encourages the model to learn more uncertain, hence generalizable, features.
@@
CMC (Contrastive Multiview Coding) regularizes learning by enforcing invariance to different views of the same scene. This invariance is a form of regularization, similar to dropout variations that encourage the model to learn features that are invariant to specific changes in the input.
@@
SELF-LABELLING approach involves pseudo-labeling, where the network assigns labels to its inputs and learns from these self-generated labels. This process regularizes the model by forcing it to generalize from its initial, potentially noisy, self-generated labels, akin to label smoothing which also deals with label noise.
@@ 5.2.4 CIFAR-100
Machine learning algorithms must include loss functions to gauge how well a model is doing on the job at hand and acts as a feedback signal to adjust the model’s parameters during training. By identifying and quantifying the discrepancy between expected output and the actual target labels, they play a critical part in directing the learning process. Over the past few years, contrastive learning has appeared as an effective approach in the realm of unsupervised representation learning, revolutionizing how neural networks can be trained without labeled data. This approach hinges on the idea of learning meaningful representations by increasing the similarity among positive pairs and decreasing the similarity of negative pairs. Among the several contrastive loss functions, the InfoNCE loss (Noise Contrastive Estimation) stands as a common choice. It operates by introducing noise to the input data to create negative pairs, enabling the model to differentiate between real positive pairs and artificially generated negatives. Another widely used loss is the NT-Xent loss [18] (Normalized Temperature-Scaled Cross Entropy), which extends InfoNCE by incorporating a temperature parameter to scale the pairwise similarities. Through a softmax function, it computes the negative log-likelihood of positive pairs, enhancing the overall contrastive learning process. SimCLR loss (Simple Contrastive Learning) represents yet another approach, harnessing multiple augmentations of the input data to build diverse views of the same instance. It seeks to minimize the similarity between views of different instances while increasing the correspondence between varied augmented views of the same instance. For learning representations or embeddings, the triplet loss [52] has gained popularity. It utilizes an anchor, a positive example that is close to it, and a negative example that is distant. The objective of the loss function is to reduce the distance between the anchor and a positive example while increasing the distance from the negative example, effectively creating a margin between them. The concept of margin loss also focuses on pairwise comparisons, but its objective is to ensure a margin between the representations of positive and negative pairs. By creating pairs of instances and optimizing the model, the loss pushes positive pairings closer together while pushing negative pairings farther apart. In response to the shortcomings of triplet loss, the Circle Loss[53] was recently developed. By establishing a decision boundary in a hypersphere around instances and incorporating a flexible buffer based on the instance distances, it encourages positive pairs to lie within the boundary while keeping negative pairs outside. These are just a few instances of the wide array of contrastive loss functions employed in various research works. The selection of a loss function is determined by factors like the specific problem, the nature of the data, and the required characteristics of the learned representations. Researchers often experiment with multiple loss functions to determine the most suitable one for their specific task [54]. As contrastive learning progresses, more innovative loss formulations and improvements are likely to emerge, opening up new possibilities for unsupervised representation learning .

### v2

@@ 3.0.2 Knowledge Distillation Methods
Caron et al. [45] in 2021 introduced knowledge distillation into SSL to enhance the training of ViTs. Their method, DINO, comprises two identical networks termed student and teacher networks, each with distinct parameters. Both networks include an encoder or base network, utilizing either ViTs or ResNet50, along with a projection head placed on top of the encoder network. The DINO approach involves generating multiple crops from two perspectives: a local perspective and a global perspective. The global crop perspective is fed to the teacher network, while the local crop perspective is fed to the student network. The local crop is a subset of the global perspective, exhibiting overlap between the two. The teacher network guides updates in the student network, but gradient stop is applied to halt backpropagation in the teacher network. Meanwhile, the teacher network weights are adjusted using the EMA (exponential moving average) of the weights from student network.
Li et al. (2022) [12] proposed Efficient Self-Supervised ViTs (EsViT) to achieve a optimized balance between accuracy and efficiency, reducing the cost in building state-of-the-art SSL vision systems while demonstrating enhanced scaling performance on accuracy versus throughput and model size. They introduced a multi-stage Transformer architecture, combining networks to address the issue of semantic information loss when transitioning from a single-stage transformer to a multi-stage architecture. Employing a distillation learning strategy targeting correspondence between patches, they introduced a new non-contrastive region-matching pre-training task, enhancing the model’s ability to learn intricate regional dependencies and substantially improve the quality of visual representations learned.
@@
Xie et al. (2021) [67] introduced MoBY, a SSL framework tailored for ViTs. This approach combines MoCo v2 and BYOL approaches, yielding notably high accuracy on the ImageNet-1K dataset. MoBY adopts the key queue, momentum design, and contrastive loss from MoCo v2, and it also take asymmetric data augmentations, asymmetric encoders, and momentum scheduler from BYOL. Employing a knowledge distillation approach, MoBY employees a pair of neural networks (online and target) to enhance mutual learning. Both encoders includes a backbone and a projector head, with the online encoder featuring an additional prediction head, rendering the two encoders non-uniform. Gradients progressively optimizes the online encoder, while the target encoder is a moving average of the online encoder updated through momentum. MoBY is evaluated across diverse downstream tasks, such as object detection and semantic segmentation, showcasing robustness and transferability across varied visual tasks. When utilizing DeiT-S and Swin-Transformers as backbone architectures, MoBY achieves top-1 accuracy rates of 72.8% and 75.0%, respectively, on ImageNet-1K.
@@ 3.0.3 Clustering Methods
Su et al. (2023) [70] introduced a clustering-based SSL framework tailored for ViTs, aiming to enhance performance of tasks related to object detection and segmentation. Their proposed methodology, Feature-Level SSL (FLSL), works at two levels of semantics: intra-view, which deals with individual image, and inter-view which considers over an entire dataset. FLSL utilizes a bi-level feature clustering approach, integrating mean-shift clustering intrinsic to transformers for extracting modes as patterns with a k-means-based SSL technique, ensuring semantic coherence of extracted representations both locally and globally.
@@ 3.0.4 Generative Methods
Chen et al. (2020) introduced iGPT, an image-based version of GPT, for self-supervised visual learning. Unlike ViT, which uses patch embedding, iGPT directly resizes and flattens images into lower-resolution sequences, which are then fed into a GPT-2 [72] model for autoregressive pixel prediction. Despite its significant computational cost, iGPT achieves high accuracy on CIFAR-10, surpassing supervised Wide ResNet.
Bao et al. (2022) [73] introduced BEIT, that is Bidirectional Encoder representation from Image Transformers is a self-supervised vision representation model. Instead of pixel-wise generation, they adopted MIM as a task to pretrain ViTs in a self-supervised manner. BEIT, grounded on a BERT-style visual Transformer, reconstructs masked images in the latent space. The approach involves randomly masking parts of image pieces and predicting the visual tokens relevant to these hidden pieces. The main motivation behind this pre-training technique is to restore the base visual tokens utilizing corrupted image patches. The standard visual Transformer serves as the backbone network. Evaluation using thorough fine-tuning for tasks such as image classification and semantic segmentation demonstrated competitive performance without human annotation.
He et al., (2022) [53] proposed MAE (Masked Autoencoders), in this methodology parts of input images are randomly masked and subsequently reconstructed the pixels that were masked, enabling optimized training of large models that demonstrate strong generalize across downstream tasks. This method utilizes an asymmetric encoder-decoder setup, where the encoder only looks on the visible parts of image, and a small decoder rebuilds the original image usingthe learned representation and mask tokens. By hiding a large proportion of the input image, e.g., 75%, MAE creates a challenging and useful self-supervised learning task. This scalable method allows large models to train efficiently and effectively, resulting in powerful models that generalize well. For instance, a standard ViT-Huge model achieved the highest accuracy of (87.8%) among techniques using only ImageNet-1K data. Supervised pre-training was outperformed by transfer learning in downstream task, showing strong potential for scaling behavior.
@@ 3.0.5 Hybrid Methods
Quan et al. (2023) [74] introduced a novel SSL model, based on the global contrast-masked autoencoder (GCMAE), which adeptly extracts both global and local features from pathological images. GCMAE leverages masking image reconstruction and contrastive learning as self-supervised pretext tasks, enabling the encoder to proficiently depict local-global features. This model comprises of an initializer, an encoder, a tile feature extractor, and a global feature extractor facilitating image reconstruction and contrastive learning tasks. In this investigation, an asymmetric encoder-decoder architecture is employed. ViT serves as the backbone for the encoder, pretrained using MAE. Its module related to patches feature extraction is comprised of eight transformer blocks, forming an asymmetric structure in conjunction with an encoder featuring of about 12 transformer blocks (ViT-base). This asymmetry facilitates decoupling of the encoder and decoder, fostering the acquisition of more generalized representations. Extraction of global feature is achieved via utilizing contrastive learning. The hidden representation Vpatch(vis) is used to reconstruct image through decoder and update a memory bank (B). This updation of B is executed to store the global features utilizing momentum coefficient t. The dynamically update is conducted to pursue the feature vectors of image data using memory bank as a fixed size queue. The various features that are randomly picked from the memory bank are used as a negative samples that helps to reduce the effects of the constraint caused by batch size on the heavy performance. Notably, the momentum adjustment mechanism is designed distinctively from MoCo’s model parameter update, it employ an encoder and a memory bank to facilitate contrastive learning. Vpath(vis) is depicted as the encoders output in the current epoch in the latent representation, where as the vpatch(vis) represented the previous epoch latent representation stored in B. Particularly, the latent demonstrations at the output of encoder as vpatch(vis) in the current epoch, The cost function combines both, the weighted sum of MSE loss for tile feature extraction and the NCE loss for global feature extraction. This formulation aims to minimize the gap between similar features, enhancing the model’s generalization and accuracy, particularly in cross-dataset transfer learning tasks.
@@
Fang et al. (2023) [75] proposed a hybrid SSL framework (HSL) for Hyperspectral Image (HSI) classification. Leveraging MIM and contrastive learning to enhance its effectiveness. Employing a ViT as the backbone network within an asymmetric encoder-decoder architecture with two branches, HSL achieve improved HSI classification utilizing contrastive pre-training and self-supervised generative tasks by effectively extracting spatial and spectral details. To facilitate image reconstruction through masked inputs, each branch connects intermediate features of the encoder to the decoder at the corresponding level. Contrastive learning is accomplished by utilizing the intermediate features that are obtained from two encoders that share parameters across both branches. Utilizing standard data augmentation methods, HSL generates several views of input images. Then each image view undergoes random masking technique such as MAE. Initially, image segments with mask are reorganised into a series of patches and later processed by HSL encoder and decoder to extract patterns. Extracted features are then pass through an adaptive average pooling layer and a projection head to achieve the contrastive learning task of capturing image resemblance. The decoder, which also includes multiple Transformer encoders, rebuilds patches after certain patches have been masked. Moreover, the Transformer Encoder retains the initial structure while the two output features from the encoder are used to compute contrastive loss. The loss is calculated using the decoded sequence produced by the decoder and the sequence that precedes the encoder masking, efficiently completing the masked image rebulding process. In summary, HSL combines contrastive pre-training and self-supervised generative tasks, facilitating effectiveness for downstream HSI classification task.
@@ 4 Regularization Techniques
Bootstrap Your Own Latent (BYOL) [28] employs a teacher-student framework, where the student network predicts the output of a teacher network on augmented inputs. The teacher network is updated as a moving average of the student’s parameters, ensuring learning consistency. This method mirrors the stabilization effect seen in regularization techniques like batch normalization and parameter sharing.
@@
Momentum Contrast (MoCo) [21] leverages a momentum-based moving average of the query encoder, establishing a large and consistent dictionary. This approach acts as a dynamic regularization mechanism, ensuring stability in the representation of dictionary keys over time. This stability is vital for the extraction of stable and generalizable features, akin to how traditional regularization methods prevent overfitting by maintaining consistency in model parameters.
@@
SimCLR [22] utilizes a straightforward contrastive learning framework, where its regularization is primarily through aggressive data augmentation techniques. These augmentations include random cropping, resizing, color distortion, and Gaussian blur, enabling the model to learn invariant and robust features. This is similar to how techniques like dropout and data augmentation in traditional settings prevent over-reliance on specific input features.
@@
DINO (Self-Distillation with No Labels) [45] utilizes self-distillation for regularization, where the student network mimics the teacher network, updated as an exponential moving average of the student’s parameters. This method is reminiscent of techniques like early stopping and weight decay, which aim to prevent overfitting by smoothing the learning trajectory.
@@
Barlow Twins [49] method aims to align the cross-correlation matrix between the outputs of twin networks with the identity matrix. This alignment encourages the model to learn features invariant to input distortions, functioning as a regularization technique by discouraging the learning of redundant features.
@@
SwAV (Swapping Assignments between Views) [27] regularization is unique, involving clustering approaches and consistency enforcement between cluster assignments for different views of the same image. This approach is akin to enforcing a constraint (regularization) on the feature space to enhance generalizability.
@@
TiCo (Time-Contrastive Consistent Learning of Visual Representations) [76] focuses on learning temporally consistent representations, acting as a regularization mechanism by ensuring robustness to temporal variations in data. This is similar to entropy regularization, which encourages the model to learn more uncertain, hence generalizable, features.
@@
CMC (Contrastive Multiview Coding) [77] regularizes learning by enforcing invariance to different views of the same scene. This invariance is a form of regularization, similar to dropout variations that encourage the model to learn features that are invariant to specific changes in the input.
@@
SELF-LABELLING [69] approach involves pseudo-labeling, where the network assigns labels to its inputs and learns from these self-generated labels. This process regularizes the model by forcing it to generalize from its initial, potentially noisy, self-generated labels, akin to label smoothing which also deals with label noise.
@@ 5.3 Loss functions
Machine learning algorithms must include loss functions to gauge how well a model is doing on the job at hand and acts as a feedback signal to adjust the model’s parameters during training. By identifying and quantifying the discrepancy between expected output and the actual target labels, they play a critical part in directing the learning process. Over the past few years, contrastive learning has appeared as an effective approach in the realm of unsupervised representation learning, revolutionizing how neural networks can be trained without labeled data. This approach hinges on the idea of learning meaningful representations by increasing the similarity among positive pairs and decreasing the similarity of negative pairs. Among the several contrastive loss functions, the InfoNCE loss [8] (Noise Contrastive Estimation) stands as a common choice. It operates by introducing noise to the input data to create negative pairs, enabling the model to differentiate between real positive pairs and artificially generated negatives. Another widely used loss is the NT-Xent loss [22] (Normalized Temperature-Scaled Cross Entropy), which extends InfoNCE by incorporating a temperature parameter to scale the pairwise similarities. Through a softmax function, it computes the negative log-likelihood of positive pairs, enhancing the overall contrastive learning process. SimCLR loss (Simple Contrastive Learning) represents yet another approach, harnessing multiple augmentations of the input data to build diverse views of the same instance. It seeks to minimize the similarity between views of different instances while increasing the correspondence between varied augmented views of the same instance. For learning representations or embeddings, the triplet loss [81] has gained popularity. It utilizes an anchor, a positive example that is close to it, and a negative example that is distant. The objective of the loss function is to reduce the distance between the anchor and a positive example while increasing the distance from the negative example, effectively creating a margin between them. The concept of margin loss also focuses on pairwise comparisons, but its objective is to ensure a margin between the representations of positive and negative pairs. By creating pairs of instances and optimizing the model, the loss pushes positive pairings closer together while pushing negative pairings farther apart. In response to the shortcomings of triplet loss, the Circle Loss[82] was recently developed. By establishing a decision boundary in a hypersphere around instances and incorporating a flexible buffer based on the instance distances, it encourages positive pairs to lie within the boundary while keeping negative pairs outside. These are just a few instances of the wide array of contrastive loss functions employed in various research works. The selection of a loss function is determined by factors like the specific problem, the nature of the data, and the required characteristics of the learned representations. Researchers often experiment with multiple loss functions to determine the most suitable one for their specific task [83]. As contrastive learning progresses, more innovative loss formulations and improvements are likely to emerge, opening up new possibilities for unsupervised representation learning .